### v2

# A lower bound on the number of edges in DP-critical graphs

Abstract.
A graph $G$ is $k$-critical (list $k$-critical, DP $k$-critical) if $\chi(G)=k$ ($\chi_{\ell}(G)=k$, $\chi_{\mathrm{DP}}(G)=k$) and for every proper subgraph $G^{\prime}$ of $G$, $\chi(G^{\prime})<k$ ($\chi_{\ell}(G^{\prime})<k$, $\chi_{\mathrm{DP}}(G^{\prime})<k$). Let $f(n,k)$ ($f_{\ell}(n,k),f_{\mathrm{DP}}(n,k)$) denote the minimum number of edges in an $n$-vertex $k$-critical (list $k$-critical, DP $k$-critical) graph. Our main result is that if $k\geq 5$ and $n\geq k+2$, then
$f_{\mathrm{DP}}(n,k)>\left(k-1+\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil^{-1}\right)\frac{n}{2}.$
This is the first bound on $f_{\mathrm{DP}}(n,k)$ that is asymptotically better than the well-known bound on $f(n,k)$ by Gallai from 1963. The result also yields a slightly better bound on $f_{\ell}(n,k)$ than the ones known before.
Mathematics Subject Classification: 05C07, 05C15, 05C35.
Keywords: Color-critical graphs, DP-coloring, sparse graphs.
Peter Bradshaw received support from NSF RTG grant DMS-1937241 and an AMS Simons Travel Grant.
Ilkyoo Choi is supported by the Hankuk University of Foreign Studies Research Fund, the Institute for Basic Science (IBS-R029-C1), and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-23324220).
Alexandr Kostochka is supported in part by NSF Grant DMS-2153507 and by NSF RTG Grant DMS-1937241.
Jingwei Xu is supported in part by Campus Research Board Award RB24000 of the University of Illinois Urbana-Champaign.
1. Introduction
In this paper, a graph has neither parallel edges nor loops. A multigraph may have parallel edges, but no loops. Given a multigraph $G$, let $V(G)$ and $E(G)$ denote the vertex set and edge set, respectively, of $G$. Given a vertex $v\in V(G)$, the degree of $v$, denoted $d_{G}(v)$, is the number of edges incident with $v$. The neighborhood $N_{G}(v)$ of a vertex $v$ is the set of vertices adjacent to $v$. Note that $|N_{G}(v)|\leq d_{G}(v)$, and equality holds if and only if $v$ has no incident parallel edges. A vertex of degree $d$ (at least $d$) is a $d$-vertex ($d^{+}$-vertex). For vertex subsets $S_{1}$ and $S_{2}$, let $E_{G}(S_{1},S_{2})$ denote the set of edges $xy\in E(G)$ where $x\in S_{1}$ and $y\in S_{2}$. For integers $a,b$ with $a\leq b$, let $[a,b]$ denote the set $\{a,a+1,\ldots,b\}$ of integers.
1.1. Known results on proper coloring
A (proper) $k$-coloring of a graph $G$ is a mapping $\varphi\,:\,V(G)\to[1,k]$ such that $\varphi(u)\neq\varphi(v)$ for each $uv\in E(G)$. The chromatic number of $G$, denoted $\chi(G)$, is the minimum positive integer $k$ for which $G$ has a proper $k$-coloring. A graph $G$ is $k$-colorable if $\chi(G)\leq k$. For a positive integer $k$, a graph $G$ is $k$-critical if $\chi(G)=k$, but every proper subgraph of $G$ is (k1)(k-1)-colorable.
The notion of $k$-critical graphs was introduced and systematically studied by Dirac [8, 9, 10, 12, 11, 13] starting from 1951. In particular, Dirac considered the minimum number $f(n,k)$ of edges in an $n$-vertex $k$-critical graph. With this notation, $f(k,k)=\binom{k}{2}$ and $f(k+1,k)$ is not well defined because there is no $(k+1)$-vertex $k$-critical graph. Since every vertex of a $k$-critical graph has degree at least k1k-1, we trivially have $f(n,k)\geq(k-1)\frac{n}{2}$. Even this simple bound yields the Heawood Formula [19] that every graph $G$ embeddable in an orientable surface $S_{\gamma}$ with $\gamma\geq 1$ handles satisfies $\chi(G)\leq\lfloor c_{\gamma}\rfloor$, where $c_{\gamma}=\frac{7+\sqrt{1+48\gamma}}{2}$. The well-known Brooks’ Theorem [5] implies that for $k\geq 4$ and $n\geq k+2$, $f(n,k)\geq\frac{(k-1)n+1}{2}$. Dirac sharpened Heawood’s result by showing that for $\gamma\geq 1$, every graph embeddable in $S_{\gamma}$ with chromatic number $\lfloor c_{\gamma}\rfloor$ contains the complete graph with $\lfloor c_{\gamma}\rfloor$ vertices as a subgraph. For this, he used the following lower bound on $f(n,k)$:
Theorem 1.1 (Dirac [11]).
If $k\geq 4$ and $n\geq k+2$, then
(1) $f(n,k)\geq\left(k-1\right)\frac{n}{2}+\frac{k-3}{2}.$
Fifteen years later, Dirac [13] described the $k$-critical graphs for which the bound (1) is exact.
In his fundamental papers [16] and [17] from 1963, Gallai proved a series of important properties of color-critical graphs. In particular, he found the exact values of $f(n,k)$ for all $n\in[k+2,2k-1]$.
Recall that a block in a graph $G$ is a maximal connected subgraph of $G$ that has no cut vertices. A Gallai tree is a connected graph whose every block is a complete graph or an odd cycle, and a Gallai forest is a graph in which every connected component is a Gallai tree.
Theorem 1.2 (Gallai [17]).
Let $k\geq 4$, and let $G$ be a $k$-critical graph. If $B$ is the set of (k1)(k-1)-vertices in $G$, then $G[B]$ is a Gallai forest.
Gallai [16] showed that his Theorem 1.2 implies the following lower bound on $f(n,k)$: if $k\geq 4$ and $n\geq k+2$, then
(2) $f(n,k)\geq\left({k-1}+\frac{k-3}{k^{2}-3}\right)\frac{n}{2}.$
For large $n$, this bound is much stronger than (1).
Krivelevich [27, 28] improved the lower bound on $f(n,k)$ in (2) to
(3) $f(n,k)\geq\left(k-1+\frac{k-3}{k^{2}-2k-1}\right)\frac{n}{2}$
and demonstrated nice applications of his bound. Then, Kostochka and Stiebitz [24] proved that for $k\geq 6$ and $n\geq k+2$,
(4) $f(n,k)\geq\left(k-1+\frac{2(k-3)}{k^{2}+6k-9-6/(k-2)}\right)\frac{n}{2}.$
Note that the last term in the bound (4) is asymptotically (in $k$) twice larger than the one in (2). Then, Kostochka and Yancey [25] proved an asymptotically exact bound: if $k\geq 4$ and $n\geq k,\,n\neq k+1$, then
(5) $f(n,k)\geq\left(k-1+\frac{k-3}{k-1}\right)\frac{n}{2}-\frac{k(k-3)}{2(k-1)}.$
This bound on $f(n,4)$ yields a short proof of the well-known theorem by Grötzsch [18] that every triangle-free planar graph is $3$-colorable.
1.2. Known results on list coloring
List coloring was introduced independently by Vizing [33] and Erdős, Rubin, and Taylor [15]. For a set $X$, let $\operatorname{Pow}(X)$ denote the power set of a set $X$, and denote $\bigcup_{v\in X}f(v)$ by $f(X)$. A list assignment for a graph $G$ is a function $L\colon V(G)\to\operatorname{Pow}(Y)$, where $Y$ is a set, whose elements are referred to as colors. For each $u\in V(G)$, the set $L(u)$ is called the list of $u$ and its elements are said to be available for $u$. A proper coloring $\varphi\colon V(G)\to Y$ is called an $L$-coloring if $\varphi(u)\in L(u)$ for each $u\in V(G)$. A graph $G$ with a list assignment $L$ is said to be $L$-colorable if it admits an $L$-coloring. The list chromatic number $\chi_{\ell}(G)$ of $G$ is the least positive integer $k$ such that $G$ is $L$-colorable whenever $L$ is a list assignment for $G$ with $|L(u)|\geq k$ for all $u\in V(G)$. If $L(u)=[1,k]$ for a positive integer $k$ for all $u\in V(G)$, then $G$ is $L$-colorable if and only if it is $k$-colorable; in this sense, list coloring generalizes proper coloring. In particular, $\chi_{\ell}(G)\geq\chi(G)$ for all graphs $G$.
The definition of critical graphs can be naturally extended to the list coloring setting. A graph $G$ is list $k$-critical if $\chi_{\ell}(G)=k$, but $\chi_{\ell}(G^{\prime})\leq k-1$ for each proper subgraph $G^{\prime}$ of $G$. So, we can define $f_{\ell}(n,k)$ to be the minimum number of edges in an $n$-vertex list $k$-critical graph.
A list assignment $L$ for a graph $G$ is called a degree-list assignment if $|L(u)|\geq d_{G}(u)$ for all $u\in V(G)$. A fundamental result of Borodin [3] and Erdős, Rubin, and Taylor [15] provides a complete characterization of all graphs that are not $L$-colorable with respect to some degree-list assignment $L$. This result can be viewed as a generalization of Theorem 1.2.
Theorem 1.3 (Borodin [3]; Erdős, Rubin, and Taylor [15]).
Let $G$ be a connected graph and $L$ be a degree-list assignment for $G$. If $G$ is not $L$-colorable, then $G$ is a Gallai tree; furthermore, $|L(u)|=d_{G}(u)$ for all $u\in V(G)$ and if $u,v\in V(G)$ are two adjacent non-cut vertices, then $L(u)=L(v)$.
Kostochka, Stiebitz, and Wirth [22] and independently Thomassen [32] described for each Gallai tree $G$ the degree-list assignments $L$ such that $G$ is not $L$-colorable. Kostochka and Stiebitz [23] proved that for $n\geq k+2$, Dirac’s bound (1) holds for $f_{\ell}(n,k)$.
Exactly as Theorem 1.2 implies (2), Theorem 1.3 yields the same lower bound on $f_{\ell}(n,k)$. Kostochka and Stiebitz [24] improved (2) for $k\geq 9$: the last term in their bound is asymptotically (in $k$) $1.5$ times larger than the one in (2). Then, in a series of papers by Kierstead and Rabern [20], Cranston and Rabern [6], and Rabern [29, 30], this bound was significantly improved and also extended to $4\leq k\leq 8$. The best known bounds are due to Rabern [29, 30]:
Theorem 1.4 (Rabern [29, 30]).
For $k\geq 4$ and $n\geq k+2$,
[…]
In fact, the bounds were proved not only for $f_{\ell}(n,k)$ but also for the analogous parameter of paintability, which is defined in terms of an on-line list coloring process where available colors are revealed one at a time (see [29, 30] for a complete definition). The last term in Theorem 1.4 is asymptotically (in $k$) twice larger than the one in (2). Below is a table similar to ones in [6, 20, 29] with an added column of our results in this paper.
$k$-critical list $k$-critical list and DP $k$-critical
$k$ Ga [16] KY [25] KS [24] KR [20] CR [6] Ra [29, 30] This paper
4 03.0769 03.3333 03.1000
5 04.0909 04.5000 04.0983 04.1000 04.1176 04.1666
6 05.0909 05.6000 05.1052 05.1076 05.1214 05.1666
7 06.0869 06.6666 06.1149 06.1192 06.1296 06.1666
8 07.0819 07.7142 07.1127 07.1167 07.1260 07.1428
9 08.0769 08.7500 08.0838 08.1093 08.1130 08.1213 08.1428
10 09.0721 09.7777 09.0793 09.1054 09.1088 09.1162 09.1250
15 14.0540 14.8571 14.0610 14.0863 14.0884 14.0930 14.1000
20 19.0428 19.8947 19.0490 19.0718 19.0733 19.0762 19.0833
Table 1. Current lower bounds on the average degree of $k$-critical and list $k$-critical graphs. Best results for list $k$-critical graphs are marked with an asterisk.
1.3. Known results on DP-coloring and our results
In this paper we focus on a generalization of list coloring that was recently introduced by Dvořák and Postle [14]; they called it correspondence coloring, and we call it DP-coloring for short. Dvořák and Postle invented DP-coloring in order to approach an open problem about list colorings of planar graphs with no cycles of certain lengths. This hints at the usefulness of DP-coloring for graph coloring problems, in particular list coloring problems.
For a multigraph $G$, a (DP-)cover of $G$ is a pair $(H,L)$, where $H$ is a multigraph and $L\colon V(G)\to\operatorname{Pow}(V(H))$ is a function such that
  • The family $\{L(u):u\in V(G)\}$ forms a partition of $V(H)$.
  • For each $u\in V(G)$, $H[L(u)]$ is an independent set.
  • For each $u,v\in V(G)$, if $|E_{G}(u,v)|=s$, then $E_{H}(L(u),L(v))$ is the union of $s$ matchings (where each matching is not necessarily perfect and possibly empty).
An $(H,L)$-coloring of $G$ is a function $\phi:V(G)\rightarrow\bigcup_{v\in V(G)}L(v)$ with $\phi(v)\in L(v)$ such that $\{(v,\phi(v)):v\in V(G)\}$ is independent in $H$. In view of this definition, we often refer to the vertices of $H$ as colors. The DP chromatic number $\chi_{\mathrm{DP}}(G)$ of a multigraph $G$ is the least positive integer $k$ such that $G$ has an $(H,L)$-coloring whenever $(H,L)$ is a cover of $G$ with $|L(u)|\geq k$ for all $u\in V(G)$. Every list coloring problem can be represented as a DP-coloring problem. In particular, $\chi_{\mathrm{DP}}(G)\geq\chi_{\ell}(G)$ for all multigraphs $G$.
A cover $(H,L)$ of a multigraph $G$ is called a degree-cover if $|L(u)|\geq d_{G}(u)$ for all $u\in V(G)$. We say a multigraph $G$ is DP degree-colorable if $G$ has an $(H,L)$-coloring whenever $(H,L)$ is a degree-cover.
A multigraph $G$ is DP $k$-critical if $\chi_{\mathrm{DP}}(G)=k$ and $\chi_{\mathrm{DP}}(G^{\prime})\leq k-1$ for every proper subgraph $G^{\prime}$ of $G$. Let $f_{\mathrm{DP}}(n,k)$ denote the minimum number of edges in an $n$-vertex DP $k$-critical graph.
For a graph $G$ and a positive integer $s$, the multiple $sG$ of $G$ is the multigraph obtained from $G$ by replacing each edge $e\in E(G)$ with $s$ edges joining the endpoints of $e$. In particular, $1G=G$. A GDP-forest is a multigraph such that for every block $B$, there exist $n$ and $t$ such that $B$ is either a $tK_{n}$ or a $tC_{n}$. (A double cycle is a multigraph $2C_{n}$.) A GDP-tree is a connected GDP-forest. Note that every Gallai tree is also a GDP-tree. An example of a GDP-tree is shown in Figure 1.
Figure 1. A GDP-tree with two $K_{3}$ blocks, a $2K_{3}$ block, two $K_{2}$ blocks, a $K_{4}$ block, and a $C_{8}$ block.
Theorems 1.2 and 1.3 extend to DP-coloring as follows.
Theorem 1.5 (Dvořák and Postle [14] for graphs; Bernshteyn, Kostochka, and Pron’ [1] for multigraphs).
Let $G$ be a connected multigraph, and let $(H,L)$ be a degree-cover of $G$. If $G$ has no $(H,L)$-coloring, then $G$ is a GDP-tree.
Kim and Ozeki [21] characterized for each GDP-tree the degree-covers $(H,L)$ for which $G$ is not $(H,L)$-colorable.
Analogous to how Theorems 1.2 and 1.3 imply lower bounds on $f(n,k)$ and $f_{\ell}(n,k)$, respectively, Theorem 1.5 yields that the lower bound in (2) holds also for $f_{\mathrm{DP}}(n,k)$ when $n\geq k+2$. Note that the bound (2) does not hold for DP-coloring of general multigraphs: for example, multiples of cycles $kC_{n}$ are $2k$-regular and DP $(2k+1)$-critical.
Bernshteyn and Kostochka [2] proved that for $n\geq k+2$, Dirac’s bound (1) holds also for $f_{\mathrm{DP}}(n,k)$. The reader can find more history of the study of color-critical graphs, including seminal results by Gallai, Hajós, Sachs, and Stiebitz in the recent monographs [7, 31].
The main result of this paper is a lower bound on $f_{\mathrm{DP}}(n,k)$ that is asymptotically better than (2).
Theorem 1.6.
If $k\geq 5$, $\lambda=\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil$, and $n\geq k+2$, then
(6) $f_{\mathrm{DP}}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}.$
We do not know whether $f_{\ell}(n,k)\geq f_{\mathrm{DP}}(n,k)$ for all $n$ and $k$. However, our method for proving Theorem 1.6 also allows us to show the following bound on $f_{\ell}(n,k)$.
Theorem 1.7.
If $k\geq 5$, $\lambda=\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil$, and $n\geq k+2$, then
(7) $f_{\ell}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}.$
This is slightly better than the bound of Theorem 1.4 for $k\geq 5$, see the last column of Table 1. We do not know how to prove (7) without using DP-coloring. Thus, this is another example of usefulness of DP-coloring for proving results on list coloring. As it is, an analog of Theorem 1.6 for $k=4$ requires additional ideas to work. In [4], we use a somewhat more elaborate technique to prove that for $n\geq 11$, $f_{\mathrm{DP}}(n,4)>\frac{8}{5}n$, which is better than (6) for $k=4$.
1.4. Outline of the proof and structure of the paper
The plan to prove Theorem 1.6 is as follows.
  1. (1)
    We want to prove by induction on $n$ that if $G\neq K_{k}$ is an $n$-vertex DP $k$-critical graph with $m$ edges, then $2m\geq\left(k-1+\frac{1}{\lambda}\right)n+\frac{2}{\lambda}$, which is equivalent to
    (8) $(\lambda(k-1)+1)n-2\lambda m\leq-2.$
  2. (2)
    We wish to consider a smallest counterexample $G$ to (8) and use the minimality of $G$ to consider smaller graphs for which (8) holds. In doing so, we hope to infer some structural properties of $G$. For this, we use a weighted version of (8). If we assign $w(v)=\lambda(k-1)+1$ for each $v\in V(G)$ and $w(e)=2\lambda$ for each $e\in E(G)$, then we can rewrite (8) as
    (9) $\sum_{v\in V(G)}\left(w(v)-\sum_{e\in E(G):v\sim e}\frac{w(e)}{2}\right)\leq-2.$
    Note that vertices of degree at most k1k-1 contribute positive terms to the sum, while vertices of degree at least $k$ contribute negative terms. Therefore, in order to prove the inequality, we plan to prove that the $k^{+}$-vertices of $G$ contribute enough to the sum in order to negate the positive terms from the vertices of degree at most k1k-1.
  3. (3)
    Sometimes it is convenient to consider smaller graphs obtained from $G$ in which some vertices have lists with fewer than k1k-1 colors. For induction, it is also convenient to consider multigraphs rather than just graphs. Therefore, we develop a more general and sophisticated model in which the weights of vertices with smaller list sizes are less than $\lambda(k-1)+1$ and the weights of multiple edges are greater than $2\lambda$. In this new model, we still aim to prove (9).
  4. (4)
    In our new model, every positive term of (9) corresponds to a low vertex, that is, a vertex $v$ for which the list size $h(v)$ is equal to the degree $d_{G}(v)$. We let $\Lambda$ be the set of low vertices in $G$, a minimum counterexample to our new model. We show that the induced subgraph $G[\Lambda]$ is a GDP-forest.
  5. (5)
    Overall, roughly speaking, the proof of the weighted version of (9) has three parts:
    (a) We show that GDP-forests of maximum degree k1k-1 and without (k1)(k-1)-regular and (k2)(k-2)-regular blocks do not have too many edges.
    (b) We show that in a minimum counterexample $G$, $G[\Lambda]$ has neither (k1)(k-1)-regular nor (k2)(k-2)-regular blocks.
    (c) We use discharging to show that negative terms in (9) compensate for all positive ones, proving the inequality.
The structure of the paper is as follows. In Section 2, we introduce our model discussed above, which allows multiple edges and variable list sizes of the vertices. In Section 3 we discuss properties of GDP-forests. Properties of a minimum counterexample are shown in Section 4, and we finish the proof via the discharging method in Section 5. In Section 6 we prove the bound for $f_{\ell}(n,k)$. In conclusion, we briefly discuss possible further research directions.
2. The setup
We will prove a statement stronger than Theorem 1.6 in the setting of multigraphs with the language of potentials, and we will obtain Theorem 1.6 as a corollary.
Let $G$ be a multigraph, and let $h:V(G)\to[0,k-1]$ be a function. For each $v,x,y\in V(G)$, we define the potential $\rho_{G,h}(v)$ and $\rho_{G,h}(xy)$ to be as follows:
$\hfill\rho_{G,h}(v)=\begin{cases}h(v)\lambda+1&\textrm{if }h(v)=k-1\\ h(v)\lambda-1&\textrm{if }h(v)\in[2,k-2]\\ h(v)\lambda-2&\textrm{if }h(v)\in\{0,1\}.\end{cases}\phantom{0}\hfill\phantom{0}\rho_{G,h}(xy)=\begin{cases}0&\textrm{if }|E_{G}(x,y)|=0\\ 1-(2\lambda+1)|E_{G}(x,y)|&\textrm{otherwise.}\end{cases}\hfill$
In other words, if $x$ and $y$ are joined by a single edge in $G$, then $\rho_{G,h}(xy)=-2\lambda$, and each additional edge joining $x$ and $y$ adds $(2\lambda+1)$ to $\rho_{G,h}(xy)$.
Given a vertex subset $A\subseteq V(G)$, we define the potential of $A$ as
$\rho_{G,h}(A)=\sum_{x\in V(A)}\rho_{G,h}(x)+\sum_{xy\in\binom{A}{2}}\rho_{G,h}(xy).$
When $A$ consists of a single vertex $v$, we often write $\rho_{h}(v)=\rho_{G,h}(\{v\})$. We also write $\rho_{h}(G)=\rho_{G,h}(V(G))$. We often omit the subscripts $G$ and $h$ when they are clear from the context.
Let $\mathbb{N}$ denote the set of positive integers. Given a multigraph $G$, let $h:V(G)\to\mathbb{N}\cup\{0\}$. A cover $(H,L)$ of $G$ is an $h$-cover of $G$ if $|L(v)|\geq h(v)$ for each $v\in V(G)$. When $h\equiv t$ for some constant $t\in\mathbb{N}$, we call an $h$-cover simply a $t$-cover. We say that $G$ is DP $h$-colorable if $G$ has an $(H,L)$-coloring for every $h$-cover $(H,L)$ of $G$. We say that $G$ is (DP) $h$-minimal if $G$ is not DP $h$-colorable, but every proper subgraph $G^{\prime}$ of $G$ is DP $h|_{V(G^{\prime})}$-colorable. Given a vertex subset $A\subseteq V(G)$, we often write $h$ for the restriction $h|_{A}$ when this does not cause confusion.
We now present our main result:
Theorem 2.1.
Let $k\geq 5$. Let $G$ be a multigraph and $h:V(G)\to[0,k-1]$. If $G$ is $h$-minimal, then one of the following holds:
  1. (1)
    $G=K_{k}$ and $h(v)\in\{k-1,k-2\}$ for each $v\in V(G)$, with at most one vertex $w\in V(G)$ satisfying h(w)=k2h(w)=k-2,
  2. (2)
    $k=5$, either $G=4K_{2}$ or $G$ is a double cycle, and $h(v)=4$ for each $v\in V(G)$, or
  3. (3)
    $\rho_{h}(G)\leq-2$.
Note that when $G$ is a graph and h(v)=k1h(v)=k-1 for each $v\in V(G)$, the statement $\rho_{h}(G)\leq-2$ is equivalent to the statement $|E(G)|\geq(k-1+\frac{1}{\lambda})\frac{|V(G)|}{2}+\frac{1}{\lambda}$. Therefore, since $4K_{2}$ and double cycles are not graphs, Theorem 2.1 implies Theorem 1.6.
A multigraph is exceptional if it is either a complete graph, or $4K_{2}$ or a double cycle in the case $k=5$.
We end this section with some lemmas for the potential function. The following well-known lemma (c.f. [26, Fact 11]) shows that our potential function is submodular.
Lemma 2.2.
If $U_{1},U_{2}\subseteq V(G)$, then $\rho_{G,h}(U_{1}\cup U_{2})+\rho_{G,h}(U_{1}\cap U_{2})\leq\rho_{G,h}(U_{1})+\rho_{G,h}(U_{2})$.
Let $G^{-}$ denote a multigraph obtained from $G$ by removing any edge.
Lemma 2.3.
Let $F$ be a copy of $K_{k}$ and let $h(v)\in\{k-2,k-1\}$ for each $v\in V(F)$, where at most one vertex $v$ has h(v)=k2h(v)=k-2. For a proper induced subgraph $F^{\prime}$ of $F^{-}$, $\rho_{h}(F^{\prime})\geq\rho_{h}(F^{-})+(k-3)(\lambda-1)-2$.
Proof.
The potential of a complete graph $K_{j}$ having $j_{1}\in\{j-1,j\}$ vertices $v$ with h(v)=k1h(v)=k-1 and $j_{2}=j-j_{1}$ vertices $u$ with h(u)=k2h(u)=k-2 is
$j_{1}((k-1)\lambda+1)+j_{2}((k-2)\lambda-1)-2\lambda\binom{j}{2}=jk\lambda-j\lambda-j_{2}\lambda+j_{1}-j_{2}-\lambda j(j-1)$
$\lambda j^{2}+jk\lambda+j_{1}-j_{2}(\lambda+1).$
In order to minimize the potential of $F^{\prime}$, we assume that $F^{\prime}$ is a complete graph. If $|V(F^{\prime})|=j$, then $\rho_{h}(F^{\prime})=-\lambda j^{2}+j(\lambda k+1)$ when $j_{2}=0$ and $\rho_{h}(F^{\prime})=-\lambda j^{2}+j(\lambda k+1)-\lambda-2$ when $j_{2}=1$. In both cases, as a function of $j$, $\rho_{h}(F^{\prime})$ is a concave quadratic function and hence is minimized at $j=1$ or j=k1j=k-1. Therefore, if $j_{2}=0$, then
$\rho_{h}(F^{\prime})\geq\min\{(k-1)\lambda+1,(k-1)(\lambda+1)\}=(k-1)\lambda+1,$
and if $j_{2}=1$, then
$\rho_{h}(F^{\prime})\geq\min\{(k-2)\lambda-1,(k-2)(\lambda+1)-1\}=(k-2)\lambda-1.$
Thus, in both cases,
$\rho_{h}(F^{\prime})\geq(k-1)\lambda+1-j_{2}(\lambda+2).$
As $\rho_{h}(F^{-})=k-j_{2}(\lambda+2)+2\lambda$, we get
$\rho_{h}(F^{\prime})-\rho_{h}(F^{-})\geq(k-1)\lambda+1-k-2\lambda=(k-3)(\lambda-1)-2.$
Lemma 2.4.
Let $k=5$, and let $F$ be an exceptional multigraph such that $h(v)=4$ for each $v\in V(F)$. Then $\rho_{h}(F)\in\{5,0,-1\}$. Moreover, if $A$ is a non-empty proper subset of $V(F^{-})$, then $\rho_{F^{-}}(A)\geq\rho_{h}(F^{-})+8$.
Proof.
As $k=5$, $\lambda=6$, so, $(k-1)\lambda+1=25$ and $2\lambda=12$. If $F=K_{5}$, then $\rho_{h}(F)=5(25)-10(12)=5$. If $F=4K_{2}$, then $\rho_{h}(F)=2(25)-12-3(13)=-1$. If $F=2C_{n}$, then $\rho_{h}(F)=25n-12n-13n=0$.
Let $A$ be a nonempty proper subset of $V(F^{-})$. We consider the following cases. If $F=4K_{2}$, then $\rho_{h}(F^{-})=12$, and $\rho_{h}(A)=25$. If $F=K_{5}$, then $\rho_{h}(F^{-})=17$, and $\rho_{F^{-}}(A)\geq 25|A|-12\binom{|A|}{2}\geq 25$. If $F$ is a double cycle, then $\rho_{h}(F^{-})=13$ and $\rho_{F^{-},h}(A)\geq 25|A|-(12+13)(|A|-1)=25$.∎
3. Properties of GDP-forests
For a positive integer $q$, the $q$-blowup of a graph $G$ is the multigraph obtained by replacing each $v\in V(G)$ with an independent set $I_{v}$ of size $q$ and replacing each edge $uv\in E(G)$ with a $K_{q,q}$ joining $I_{u}$ and $I_{v}$.
Suppose that $(H,L)$ is a DP-cover of a multigraph $G$. The $q$-blowup of $(H,L)$ is a DP-cover of $qG$ obtained by replacing each vertex $u\in V(H)$ with an independent set $I_{u}$ of size $q$, replacing each edge $uw\in E(H)$ with a $K_{q,q}$ joining $I_{u}$ and $I_{w}$, and replacing each list $L(v)$ with the set $\bigcup_{u\in L(v)}I_{u}$.
The following lemma is a partial case of a result by Kim and Ozeki [21, Theorem 5], see also [31, Theorem 1.17].
Lemma 3.1.
Let $G$ be a multigraph and $h$ be a function on $V(G)$. Let $(H,L)$ be an $h$-cover of $G$.
(a) If $G=qK_{t}$ and h(v)=(t1)qh(v)=(t-1)q for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a (t1)(t-1)-cover $(H^{\prime},L^{\prime})$ of $K_{t}$ where $H^{\prime}$ forms t1t-1 disjoint copies of $K_{t}$.
(b) If $G=qC_{2t}$ and $h(v)=2q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t}$ where $H^{\prime}$ is $C_{4t}$.
(c) If $G=qC_{2t+1}$ and $h(v)=2q$ for all vertices $v$, then $G$ has no $(H,L)$-coloring if and only if $(H,L)$ is a $q$-blowup of a $2$-cover $(H^{\prime},L^{\prime})$ of $C_{2t+1}$ where $H^{\prime}$ forms two disjoint copies of $C_{2t+1}$.
For the next lemma, we need some more definitions and notation. Let ${\rm dist}_{G}(u,v)$ denote the distance from $u$ to $v$ in a multigraph $G$. The diameter, ${\rm diam}(G)$, of a connected multigraph $G$ is $\max_{u,v\in V(G)}{\rm dist}_{G}(u,v)$. A vertex $x\in V(G)$ is peripheral if there is $y\in V(G)$ such that ${\rm dist}_{G}(u,v)={\rm diam}(G)$. The block tree of a connected multigraph $G$ is the bipartite graph $H$ whose parts are the set of cut vertices and the set of the blocks of $G$, with a cut vertex $v$ adjacent to block $B$ in $H$ if and only if $v\in B$. This definition implies that the block tree of a connected multigraph $G$ is always a tree whose leaves are some blocks of $G$. Call a block of a connected multigraph $G$ peripheral if the corresponding vertex in the block tree of $G$ is peripheral.
For a multigraph $F$, let $\widetilde{F}$ denote the underlying graph of $F$, i.e. the graph from which $F$ is obtained by multiplying some edges. In other words, $\widetilde{F}$ is the graph on $V(F)$ in which $u$ and $v$ are adjacent if and only if $|E_{F}(u,v)|\geq 1$. For a vertex $v\in V(F)$, let $\widetilde{d}(v)$ denote $d_{\widetilde{F}}(v)$, the degree of $v$ in $\widetilde{F}$, which is equal to $|N_{F}(v)|$.
For a multigraph $F$, let m(F):=|E(F)||E(F~)|m(F):=|E(F)|-|E($\widetilde${F})|. So, $2m(F)=\sum_{v\in V(F)}(d_{F}(v)-\widetilde{d}_{F}(v))$.
Given a function $h:V(F)\to\mathbb{N}\cup\{0\}$ with $h(v)\geq d_{F}(v)$ for each $v\in V(F)$, define $\sigma_{h}(F):=\sum_{v\in V(F)}(h(v)-d_{F}(v))$. For a positive integer $j$, let $V_{j}(F)$ and $V^{-}_{j}(F)$ denote the sets of vertices $v\in V(F)$ with $h(v)=j$ and $h(v)<j$, respectively, in $F$.
Lemma 3.2.
Let $k\geq 5$ and $\alpha=\frac{k-2}{2k-7}$. Suppose that $T$ is a GDP-tree and $h:V(T)\to\mathbb{N}\cup\{0\}$ satisfies
(i) $3\leq h(v)\leq k-1$ for each $v\in V(T)$,
(ii) $h(v)\geq d_{T}(v)$ for each $v\in V(T)$, and
(iii) $T$ has neither (k1)(k-1)-regular nor (k2)(k-2)-regular blocks.
Define $\Phi_{k}(T)=\alpha\sigma_{h}(T)+m(T)+|V^{-}_{k-1}(T)|-|V_{k-1}(T)|$. Then
(10) $\Phi_{k}(T)>1+\alpha.$
Proof.
We use induction on the number of blocks in $T$.
Suppose first that $T$ has exactly one block. If $T$ is a single vertex $v$, then $d_{T}(v)=0$. So, if $h(v)\leq k-2$, then $V_{k-1}(T)=\emptyset$ and by (i), $\sigma_{h}(T)=h(v)-d_{T}(v)\geq 3$, implying that $\Phi_{k}(T)=\alpha(h(v)-d_{T}(v))+1\geq 3\alpha+1>1+\alpha$. Otherwise, $|V_{k-1}(T)|=1$ and
$\Phi_{k}(T)=\alpha(h(v)-d_{T}(v))-1=\alpha(k-1)-1=\alpha-1+\frac{(k-2)^{2}}{2k-7}>1+\alpha$
for $k\geq 5$, again implying (10). Suppose $T$ has two vertices, say $v_{1}$ and $v_{2}$, and $j\geq 1$ edges, so $d_{T}(v_{1})=d_{T}(v_{2})=j$, $|V(T)|=2$ and m(T)=j1m(T)=j-1. By (i)–(iii), $j\leq k-3$. We can write $\Phi_{k}(T)$ in the form $\Phi_{k}(T)=\sum_{i=1}^{2}\Phi^{\prime}_{k}(v_{i})$, where
$\Phi^{\prime}_{k}(v_{i})=\alpha(h(v_{i})-d_{T}(v_{i}))+\frac{j-1}{2}+|V_{k-1}^{-}(T)\cap\{v_{i}\}|-|V_{k-1}(T)\cap\{v_{i}\}|.$
For each $i\in\{1,2\}$, if $h(v_{i})=k-1$, then since $j\leq k-3$ and $k\geq 5$,
$\Phi^{\prime}_{k}(v_{i})$ $=$ $\alpha(k-1-j)+\frac{j-1}{2}+0-1=\alpha(k-1)-\frac{3}{2}-\frac{3j}{4k-14}$
$\geq$ $\alpha(k-1)-\frac{3}{2}-\frac{3k-9}{4k-14}=\frac{2k^{2}-15k+34}{2(2k-7)}>1.$
If $h(v_{i})<k-1$, then
$\Phi^{\prime}_{k}(v_{i})\geq\begin{cases}\displaystyle 0+\frac{j-1}{2}+1-0\geq\frac{3}{2}&\mbox{if $j\geq 2$}\\[8.61108pt] \alpha(3-j)+0+1-0>2&\mbox{if $j=1$}\end{cases}$
Thus in all cases, $\Phi_{k}(T)=\sum_{i=1}^{2}\Phi^{\prime}_{k}(v_{i})>2\geq 1+\alpha,$ as claimed.
If $T$ has more than two vertices, then $T$ is a multiple of a cycle or a complete graph. In particular, $T$ is $j$-regular for some $1\leq j\leq k-3$. Similarly to above, we can write $\Phi_{k}(T)=m(T)+\sum_{v\in V(T)}\Phi_{k}(v)$, where
$\Phi_{k}(v)=\alpha(h(v)-d_{T}(v))+|V_{k-1}^{-}(T)\cap\{v\}|-|V_{k-1}(T)\cap\{v\}|.$
For every $v\in V_{k-1}^{-}(T)$, $\Phi_{k}(v)\geq 0+1-0=1.$ If $j\leq k-4$, then for every $v\in V_{k-1}(T)$,
(11) $\Phi_{k}(v)=\alpha(k-1-j)+0-1\geq 3\alpha-1$
In this case, $\Phi_{k}(T)\geq|V(T)|\geq 3\cdot\min\{1,3\alpha-1\}>1+\alpha$, and we are done. Thus, we suppose j=k3j=k-3. Then instead of (11), we get
(12) $\Phi_{k}(v)=\alpha(k-1-j)+0-1=2\alpha-1=\frac{3}{2k-7}>0.$
So if $|V_{k-1}^{-}(T)|+m(T)\geq 2$, then $\Phi_{k}(T)>2>1+\alpha$, as claimed. Otherwise, $G$ is a (k3)(k-3)-regular graph and $|V_{k-1}^{-}(T)|\leq 1.$ It follows that $G=K_{k-2}$ and by (12),
$\Phi_{k}(T)\geq(k-2)\left(2\alpha-1\right)=(k-2)\frac{3}{2k-7}=3\alpha>1+\alpha.$
This proves the base of induction.
Now, let $b\geq 1$, and suppose that the lemma holds for GDP-trees with at most $b$ blocks, and that $T$ has $b+1\geq 2$ blocks. Choose a peripheral block $B$ in $T$ so that the cut vertex $u$ in $B$ has the minimum number of neighbors in $B$. Let s=|V(B)|1s=|V(B)|-1.
Let $T^{\prime}=T-(V(B)-u)$. Then $T^{\prime}$ is a GDP-tree satisfying (i)–(iii) and has fewer blocks than $T$. So, by induction, $\Phi_{k}(T^{\prime})>1+\alpha$. Since $T$ is a GDP-tree, $B$ is $j$-regular for some $1\leq j\leq k-3$. In particular, $d_{T}(u)-d_{T^{\prime}}(u)=j$, and the multiplicity of every existing edge in $B$ is the same, say $q$.
Suppose first that $s=1$, so that $V(B)=\{u,v\}$. If $v\in V_{k-1}^{-}(T)$, then $\Phi_{k}(T)\geq\Phi_{k}(T^{\prime})-\alpha j+(j-1)+1-0\geq\Phi_{k}(T^{\prime})$. If $v\in V_{k-1}(T)$, then since $j\leq k-3$,
$\Phi_{k}(T)-\Phi_{k}(T^{\prime})=-\alpha j+(j-1)-1+\alpha(k-1-j)=j(1-2\alpha)-2+\alpha(k-1)$
$\geq(k-3)(1-2\alpha)-2+\alpha(k-1)=-\frac{3(k-3)}{2k-7}-2+\frac{(k-1)(k-2)}{2k-7}=\frac{k^{2}-10k+25}{2k-7}\geq 0.$
On the other hand, suppose $s\geq 2$. Then
(13) $\widetilde{B}$ is $t$-regular for some $t\geq 2$, and $j=qt\geq 2q$.
For each $\in(V(B)-u)\cap V_{k-1}(T), vv$ contributes $\alpha(k-1-j)+0-1$ to $\Phi_{k}(v)$, and for each $\in(V(B)-u)\cap V_{k-1}^{-}(T), vv$ contributes at least $1$ to $\Phi_{k}(v)$. For two adjacent vertices $x,y\in V(B)$, the $q$ edges in $E_{B}(x,y)$ contribute q1q-1 to $\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ due to the increase in $m(T)$. Moreover, for each neighbor $x\in V(B)$ of $u$, the $q$ edges in $E_{G}(u,x)$ contribute an additional $\alpha q$ to $\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ due to the decrease in $h(u)-d_{T}(u)$. Note that $s\geq\frac{j}{q}$, and $B$ has at least $(s+1)q$ edges, so $m(B)\geq(s+1)(q-1)$.
If $j\leq k-4$, then by (13), $k\geq 4+j\geq 4+t\geq 6$. In this case, each vertex $v\in V(B)-u$ contributes at least $\min\{\alpha(k-1-j)-1,1\}\geq\min\{3\alpha-1,1\}\geq\alpha$ to $\Phi_{k}(T)$. Therefore,
$\Phi_{k}(T)-\Phi_{k}(T^{\prime})$ $\geq$ $s\alpha+(s+1)(q-1)-\frac{j}{q}(q\alpha)=\alpha(s-j)+(s+1)(q-1)$
$\geq$ $\alpha(s-j)+j-s+q-1=(j-s)(1-\alpha)+q-1\geq 0,$
since the third expression is nonnegative when $s\geq j$, and the last expression is nonnegative when $j\geq s$.
Otherwise, j=k3j=k-3, so each vertex $v\in V(B)-u$ contributes at least $\min\{\alpha(k-1-j)-1,1\}=\min\{2\alpha-1,1\}=2\alpha-1$ to $\Phi_{k}(T)$. If $q\geq 2$, then $j=qt\geq 4$. Then, by (13), $k=3+j\geq 7.$ If $s=2$, then $\widetilde{B}$ is a $3$-cycle, so $j$ is even and $m(B)=\frac{3j}{2}-3\geq j-1$ since $j\geq 4$. If $s\geq 3$, then $m(B)=\frac{j-t}{2}(s+1)\geq\frac{j}{4}(3+1)=j$. Hence, $m(B)\geq j-1$.
So, $m(B)-\alpha(d_{T}(u)-d_{T^{\prime}}(u))\geq j-1-\alpha j=(j-1)(1-\alpha)-\alpha$. Therefore, since $s\geq 2$ and $j\geq 3$,
$\Phi_{k}(T)-\Phi_{k}(T^{\prime})\geq(j-1)(1-\alpha)-\alpha+s(2\alpha-1)\geq 2(1-\alpha)-\alpha+4\alpha-2=\alpha>0.$
The remaining possibility is that j=k3j=k-3 and $q=1$, which means that $B=K_{k-2}$ or $k=5$ and $B$ is a cycle. In the latter case, each $v\in T^{\prime}$ contributes at least $1\cdot(3-2)=1$, since $\alpha=1,h(v)\geq 3$, and $j=2$. Thus $\Phi_{k}(T)-\Phi_{k}(T^{\prime})>0$. In the former case, we have $d_{T^{\prime}}(u)=d_{T}(u)-j\leq 2$, so write $N_{T^{\prime}}(u)=\{u_{1},u_{2}\}$ (possibly, $u_{2}=u_{1}$).
We say a tree $T_{0}$ is good if $T_{0}$ is a GDP-tree satisfying (i)–(iii) with fewer blocks than $T$ and $V(T_{0})\subseteq V(T)-(V(B)-u)$. By induction, $\Phi_{k}(T_{0})>1+\alpha$. In the remainder of the proof, we will demonstrate $\Phi_{k}(T)-\Phi_{k}(T_{0})\geq 0$, which proves the lemma. We will often use the fact that
  1. (*)
    Each $v\in(V(B)-u)\cap V_{k-1}^{-}(T)$ contributes $1$ to $|V_{k-1}^{-}(T)\cap\{v\}|$, and
    each $v\in(V(B)-u)\cap V_{k-1}(T)$ contributes $2\alpha-1$ to $\alpha(h(v)-d_{T}(v))-|V_{k-1}(T)\cap\{v\}|$,
so the vertices in V(B)uV(B)-u collectively contribute at least $(k-3)\min\{1,2\alpha-1\}=(k-3)(2\alpha-1)$ to $\Phi_{k}(T)$.
Write $T^{\prime\prime}=T^{\prime}-u$ and observe that $T^{\prime\prime}$ is a good tree. Since $B$ is peripheral, $u_{1}$ and $u_{2}$ are in the same block, say $B_{0}$. Hence, $B_{0}$ is either a $K_{2}$ or a cycle.
Case 1: Assume $u_{1}=u_{2}$, so $N_{T^{\prime}}(u)=\{u_{1}\}$. Note that $u$ contributes no less than 1-1 to $\Phi_{k}(T)$.
If $d_{T}(u)=k-2$, then $h(u_{1})-d(u_{1})$ decreases exactly by $1$. Together with (*),
$\Phi_{k}(T)-\Phi_{k}(T^{\prime\prime})\geq-\alpha-1+(k-3)(2\alpha-1)=\alpha(2k-7)-(k-2)=0.$
If $d_{T}(u)=k-1$, then $m(B_{0})=1$ and $d_{T}(u_{1})-d_{T^{\prime\prime}}(u_{1})=2$. Together with (*),
$\Phi_{k}(T)-\Phi_{k}(T^{\prime\prime})\geq-2\alpha+1-1+(k-3)(2\alpha-1)=\alpha(2k-8)-(k-3)=\frac{k-5}{2k-7}\geq 0.$
For the remaining cases, assume $u_{1}\neq u_{2}$, so $d_{T}(u)=k-1$. As $B$ is chosen so that the cut vertex of $B$ has minimum degree in $B$ over all cut vertices of peripheral blocks of $T$, $B_{0}$ is not peripheral, and thus $T$ has at least three blocks.
Case 2.1: $|V(B_{0})|\geq 4$, so $B_{0}$ is a cycle of length at least $4$, and $u_{1}u_{2}$ is not an edge. Obtain $T_{1}$ from $T^{\prime\prime}$ by adding the edge $u_{1}u_{2}$, so $T_{1}$ is a good tree. Then $u$ contributes 1-1 to $\Phi_{k}(T)-\Phi_{k}(T_{1})$. Together with (*),
$\Phi_{k}(T)-\Phi_{k}(T_{1})\geq-1+(k-3)(2\alpha-1)=\alpha(2k-6)-(k-2)>0.$
Case 2.2: $|V(B_{0})|=3$ and $d_{T}(u_{1})=2$. Obtain $T_{2}$ from $T^{\prime\prime}$ by deleting $u_{1}$, so that $T_{2}$ is a good tree. Then $d_{T}(u_{2})-d_{T_{2}}(u_{2})=2$, and $u_{1}$ contributes to $\Phi_{k}(T)-\Phi_{k}(T_{2})$ either $\alpha(k-1-2)-1$ (when $h(u_{1})=k-1$) or at least $1$ (when $h(u_{1})\leq k-2$). As in Case 2.1, $u$ contributes 1-1 to $\Phi_{k}(T)-\Phi_{k}(T_{2})$. Together with (*),
$\Phi_{k}(T)-\Phi_{k}(T_{2})\geq-2\alpha+(3\alpha-1)-1+(k-3)(2\alpha-1)=\alpha(2k-5)-(k-1)=\frac{3}{2k-7}>0.$
By the symmetry between $u_{1}$ and $u_{2}$, the last case is as follows.
Case 2.3: $|V(B_{0})|=3$ and both $u_{1}$ and $u_{2}$ are cut vertices. For $i\in\{1,2\}$, let $B_{i}$ be a block of $T$ containing $u_{i}$ and distinct from $B_{0}$. Since $B$ is peripheral, at least one of $B_{1},B_{2}$ is also peripheral, say $B_{1}$. Then by the choice of $B$, $B_{1}$ is (k3)(k-3)-regular, and so by the previous cases, $B_{1}=K_{k-2}$. In this case, $T_{3}=T-V(B)-V(B_{1})$ is a good tree. Then $d_{T}(u_{2})-d_{T_{3}}(u_{2})=2$, and each of $u,u_{1}$ contributes 1-1 to $\Phi_{k}(T)-\Phi_{k}(T_{2})$. Together with (*) for both $B$ and $B_{1}$,
$\Phi_{k}(T)-\Phi_{k}(T_{3})\geq-2\alpha-2+2(k-3)(2\alpha-1)=2\alpha(2k-7)-2(k-2)=0.$
This proves the lemma. ∎
4. Properties of a minimum counterexample $G$
Suppose that Theorem 2.1 does not hold for some $k\geq 5$. Then there are a multigraph $G$, a function $h:V(G)\to[0,k-1]$, and an $h$-cover $(H,L)$ such that
  1. (A)
    $G$ has no $(H,L)$-coloring, but every proper subgraph $G^{\prime}$ of $G$ is DP $h|_{V(G^{\prime})}$-colorable, and
  2. (B)
    none of the claims (1)–(3) of Theorem 2.1 holds for $G$ and $h$.
Among such quadruples $(G,h,H,L)$, choose one with minimum $|G|$ and fix it for the rest of the paper.
For every $uv\in E(G)$, we assume $E_{H}(L(u),L(v))$ is the union of $|E_{G}(u,v)|$ maximal matchings. We write $L(v)=\{1_{v},2_{v},\dots,t_{v}\}$, where $t=|L(v)|$.
4.1. General observations
We make some observations about our minimum counterexample $G$ and the potentials of its subgraphs.
Observation 4.1.
$G$ has no vertex $v$ for which $h(v)=0$.
Proof.
If $V(G)=\{v\}$, then $\rho_{h}(G)=\rho_{h}(v)=-2$, so $G$ is not a counterexample. If $|V(G)|\geq 2$, then $G[\{v\}]$ is a proper subgraph of $G$ that is not DP $h$-colorable, contradicting the assumption that $G$ is $h$-minimal. ∎
Observation 4.2.
For each induced subgraph $\subseteq G, XX$ is not DP $(h-d_{G}+d_{X})$-colorable.
Proof.
As $G$ is $h$-minimal, GV(X)G-V(X) has an $(H,L)$-coloring $\varphi$. For each $x\in V(X)$, let $L^{\prime}(x)\subseteq L(x)$ consist of the colors of $L(x)$ that are not adjacent to a color of $\varphi$. Then, let $H^{\prime}$ be the subgraph of $H$ induced by the set $\bigcup_{x\in V(X)}L^{\prime}(x)$. As $G$ has no $(H,L)$-coloring, $\varphi$ cannot be extended to $X$; therefore, $X$ has no $(H^{\prime},L^{\prime})$-coloring. As $|L^{\prime}(x)|\geq h(x)-d_{G}(x)+d_{X}(x)$ for each $x\in V(X)$, $X$ is not DP $(h-d_{G}+d_{X})$-colorable. ∎
Observation 4.3.
For each vertex $v\in V(G)$, $h(v)\leq d(v)$.
Proof.
We let $X=G[v]$. By Observation 4.2, $X$ is not DP $(h-d_{G}+d_{X})$-colorable. This implies that $h(v)-d_{G}(v)+d_{X}(v)=h(v)-d(v)\leq 0$, completing the proof. ∎
We say that a vertex $v\in V(G)$ is low if $h(v)=d(v)$.
Lemma 4.4.
If $U\subseteq V(G)$ is a set of low vertices, then each block of $G[U]$ is isomorphic to $sK_{t}$ or $sC_{t}$, for some $t$ and $s$.
Proof.
Suppose that $G[U]$ contains a block $B$ that is isomorphic to neither $sK_{t}$ nor $sC_{t}$. Let $(H,L)$ be an $h$-cover of $G$ for which $G$ has no $(H,L)$-coloring. As $G$ is $h$-minimal, GV(B)G-V(B) has an $(H,L)$-coloring $\varphi$. Now, for each $v\in V(B)$, let $L^{\prime}(v)\subseteq L(v)$ consist of the colors in $L(v)$ with no neighbor in $\varphi$. Each $v\in V(B)$ is low, so $|L^{\prime}(v)|\geq d_{B}(v)$. Then, $B$ has an $(H,L^{\prime})$-coloring by Theorem 1.5, and thus $G$ is $(H,L)$-colorable, a contradiction. ∎
Lemma 4.5.
For each $j\geq 1$, if $S\subsetneq V(G)$ satisfies $|E_{G}(S,\overline{S})|\geq j$, then $\rho_{G,h}(S)\geq j(\lambda-2)+1$.
Proof.
Suppose that the lemma is false, and let $j$ be the smallest value for which the lemma does not hold. Then, $G$ has a set $S\subsetneq V(G)$ for which $\rho_{G,h}(S)\leq j(\lambda-2)$ and $|E_{G}(S,\overline{S})|\geq j$. We choose $S$ to be a counterexample to the lemma with largest size.
Let $G^{\prime}=G-S$. For each vertex $v\in V(G^{\prime})$, let $h^{\prime}(v)=\max\{0,h(v)-d_{G}(v)+d_{G^{\prime}}(v)\}=\max\{0,h(v)-|E_{G}(v,S)|\}$. By Observation 4.2, $G^{\prime}$ is not DP $h^{\prime}$-colorable. Therefore, there exists $U\subseteq V(G^{\prime})$ for which $G^{\prime}[U]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal, it follows that $(G,h)$ and $(G^{\prime},h^{\prime})$ do not agree on $U$; therefore, $U$ contains a neighbor $u$ of $S$. As $h^{\prime}(u)<k-1$, and as $G$ contains no $K_{k}$ subgraph, $\rho_{G^{\prime},h^{\prime}}(U)\leq-2$.
Now, consider the set $S^{\prime}:=U\cup S$, and write $\ell=|E_{G}(S,U)|$. As $U$ contains a neighbor of $S$, $\ell\geq 1$. As $\rho_{h}(v)-\rho_{h^{\prime}}(v)\leq(\lambda+2)|E_{G}(v,S)|$ for each $v\in U$, $\rho_{G,h}(U)\leq-2+\ell(\lambda+2)$. We have
(14) $\rho_{G,h}(S^{\prime})\leq\rho_{G,h}(U)+\rho_{G,h}(S)-2\lambda\ell\leq-2+\ell(\lambda+2)-\ell(2\lambda)+\rho_{G,h}(S)=-2-\ell(\lambda-2)+\rho_{G,h}(S).$
Now, suppose $j=1$. Then, $\rho_{G,h}(S)\leq\lambda-2$, so
$\rho_{G,h}(S^{\prime})\leq-2-\ell(\lambda-2)+\lambda-2=-2+(1-\ell)(\lambda-2)\leq-2.$
Thus, it follows from the maximality of $S$ that $S^{\prime}=V(G)$. Therefore, $\rho_{h}(G)\leq-2$, and $G$ is not a counterexample to Theorem 2.1, a contradiction. This completes the case that $j=1$.
Next, suppose that $j\geq 2$. Then, as $\rho_{G,h}(S)\leq j(\lambda-2)$, (14) implies that
(15) $\rho_{G,h}(S^{\prime})\leq-2-\ell(\lambda-2)+j(\lambda-2)=-2+(j-\ell)(\lambda-2).$
If $S^{\prime}=V(G)$, then (15) implies that $\rho_{h}(G)\leq-2$, and $G$ is not a counterexample to Theorem 2.1. Otherwise, $|E_{G}(S^{\prime},\overline{S^{\prime}})|\geq 1$, so the $j=1$ case implies that $\rho_{G,h}(S^{\prime})\geq\lambda-1$. In both cases by (15), $\ell<j$, so that there is at least one edge in $E_{G}(S,V(G^{\prime})\setminus U)$.
Now, since $|E_{G}(S,U)|=\ell$ and $|E_{G}(S,\overline{S})|\geq j$, it follows that $|E_{G}(S^{\prime},\overline{S^{\prime}})|\geq|E_{G}(S,V(G^{\prime})\setminus U)|\geq j$-$\ell$. As $1\leq j-\ell\leq j-1$, the minimality of $j$ tells us that $\rho_{G,h}(S^{\prime})>(j-\ell)(\lambda-2)$, contradicting (15). ∎
We can make several observations from Lemma 4.5.
Observation 4.6.
Each vertex subset $U\subseteq V(G)$ satisfies $\rho_{h}(U)\geq-1$.
Proof.
If $1\leq|U|\leq|V(G)|-1$, then $\rho_{h}(U)\geq\lambda-1$ by Lemma 4.5. If $|U|=|V(G)|$, then $\rho_{h}(U)=\rho_{h}(G)\geq-1$ by our assumption that $G$ is a counterexample to Theorem 2.1. ∎
Observation 4.7.
For every nonempty proper subset $S\subsetneq V(G)$, $\rho_{G}(S)\geq\lambda-1$.
Proof.
As $S\subsetneq V(G)$ and $G$ is connected, $|E_{G}(S,\overline{S})|\geq 1$, so the statement follows by setting $j=1$ in Lemma 4.5. ∎
Lemma 4.8.
If $G$ has a nonempty proper subset $S\subsetneq V(G)$ for which $|E_{G}(S,\overline{S})|=1$, then $\rho_{G,h}(S)\in\{\lambda-1,\lambda\}$.
Proof.
Suppose that $E_{G}(S,\overline{S})=\{x_{0}y_{0}\}$, where $x_{0}\in S$ and $y_{0}\in\overline{S}$. If we define $h^{\prime}:V(S)\to\mathbb{N}\cup\{0\}$ so that $h^{\prime}(x_{0})=h(x_{0})-1$ and $h^{\prime}(v)=h(v)$ for each $v\in S\setminus\{x_{0}\}$, Observation 4.2 implies that $G[S]$ is not $h^{\prime}$-colorable, and hence $S$ has a subset $X$ for which $G[X]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal, $h$ and $h^{\prime}$ do not agree on $X$, so $x_{0}\in X$. As $h^{\prime}(x_{0})<k-1$, and as $G$ contains no $K_{k}$ subgraph, $\rho_{G,h^{\prime}}(X)\leq-2$. As $\rho_{h}(x_{0})-\rho_{h^{\prime}}(x_{0})\leq\lambda+2$, we have $\rho_{G,h}(X)\leq\rho_{G,h^{\prime}}(X)+(\lambda+2)\leq-2+(\lambda+2)=\lambda$. By a symmetric argument, $\overline{S}$ has a vertex subset $Y$ containing $y_{0}$ for which $\rho_{G,h}(Y)\leq\lambda$.
Now, $\rho_{G,h}(X\cup Y)\leq\lambda+\lambda-2\lambda=0$. Therefore, Observation 4.7 implies that $V(G)=X\cup Y$, further implying $X=S$ and $\rho_{G,h}(S)\leq\lambda$. As Observation 4.7 implies that $\rho_{G,h}(S)\geq\lambda-1$, the proof is complete. ∎
Lemma 4.9.
$G$ has no vertex $v$ satisfying $h(v)=1$.
Proof.
If $h(v)=1$, then $\rho_{h}(v)=\lambda-2$, contradicting Observation 4.7. ∎
Lemma 4.10.
If $h(v)\in[2,\ldots,\frac{\lambda-1}{2}]$, then $v$ is low.
Proof.
We have $\rho_{h}(v)=h(v)\lambda-1\leq\lceil\frac{h(v)\lambda-1}{\lambda-2}\rceil(\lambda-2)$. Then, Lemma 4.5 implies that
$d(v)\leq\left\lceil\frac{h(v)\lambda-1}{\lambda-2}\right\rceil-1=h(v)+\left\lceil\frac{2h(v)-1}{\lambda-2}\right\rceil-1=h(v).$
Lemma 4.11.
$G$ has no vertex $v$ satisfying $h(v)=2$.
Proof.
Suppose that $v\in V(G)$ satisfies $h(v)=2$. By Lemma 4.10, $d(v)=2$. Suppose first that $v$ is joined to a neighbor $w$ by a pair of parallel edges. Then, $N(v)=\{w\}$.
Now, let $G^{\prime}=G-v$. Choose a color $c\in L(v)$, and obtain $H^{\prime}$ from $H$ by deleting $L(v)\cup N_{H}(c)$. We may assume without loss of generality that $|L(w)\cap N_{H}(c)|\geq 2$. Let $L^{\prime}(w)=L(w)\cap V(H^{\prime})$, and let $L^{\prime}(x)=L(x)$ for each $x\in V(G^{\prime})-w$. If $G^{\prime}$ has an $(H^{\prime},L^{\prime})$-coloring $\varphi^{\prime}$, then we can extend it to $G$ by letting $\varphi^{\prime}(v)=c$. So assume $G^{\prime}$ has no $(H^{\prime},L^{\prime})$-coloring. Then there exists a subset $U^{\prime}\subseteq V(G^{\prime})$ for which $G^{\prime}[U^{\prime}]$ has a spanning $h^{\prime}$-minimal subgraph. As $G$ is $h$-minimal and $U^{\prime}\subsetneq V(G)$, it follows that $h^{\prime}$ and $h$ do not agree on some vertex of $U^{\prime}$; therefore, $w\in U^{\prime}$. As $h^{\prime}(w)\leq k-3$, it follows from the minimality of $G$ that $\rho_{G^{\prime},h^{\prime}}(U^{\prime})\leq-2$, so that $\rho_{G,h}(U^{\prime})\leq-2+\lambda+(\lambda+2)=2\lambda$. Then,
$\rho_{G,h}(U^{\prime}+v)\leq 2\lambda-(4\lambda+1)+(2\lambda-1)=-2,$
contradicting Observation 4.6.
Otherwise, suppose $N(v)=\{w,u\}$, where $u\neq w$. By Lemma 4.9, $|L(u)|\geq 2$ and $|L(w)|\geq 2$. We may assume that $\{1_{v}1_{u},2_{v}2_{u},1_{v}1_{w},2_{v}2_{w}\}\subseteq E(H)$.
Obtain $G^{\prime}$ from GvG-v by adding an edge $uw$. Also, let $(H^{\prime},L)$ be an $h$-cover of $G^{\prime}$ obtained by deleting $L(v)$ and then adding the matching $\{1_{u}2_{w},2_{u}1_{w}\}$. (Here, we consider $h$ restricted to $V(G^{\prime})$.) If $G^{\prime}$ has an $(H^{\prime},L)$-coloring $\varphi$, then we can extend $\varphi$ to $v$, because if $\varphi(u)$ forbids the color $i_{v}$ for $v$, then (due to our new edge $uw$) $\varphi(w)$ does not forbid the color $(3-i)_{v}$ for $v$. Thus, $G^{\prime}$ has an $h$-minimal subgraph $G^{\prime\prime}$ containing $u$ and $w$. Let $U=V(G^{\prime\prime})$.
As $|V(G^{\prime})|<|V(G)|$, it follows from our induction hypothesis that $\rho_{G^{\prime},h}(U)\leq-2$ or $G^{\prime\prime}$ is exceptional. We consider several cases.
First, suppose that $k=5$, $h(x)=4$ for each $x\in U$, and that $G^{\prime\prime}$ is a copy of $4K_{2}$ or $2C_{t}$ for some integer $t$. Since each adjacent pair of vertices in $G^{\prime\prime}$ has edge multiplicity at least $2$, and as $u,w\in U$, we have $uw\in E(G)$. Furthermore, as $G^{\prime\prime}$ is exceptional, $\rho_{G^{\prime\prime},h}(U)\leq 0$. Therefore,
$\rho_{G,h}(U\cup\{v\})\leq 0+(2\lambda+1)-2(2\lambda)+(2\lambda-1)=0.$
Thus, Observation 4.7 implies that $V(G)=U\cup\{v\}$. Furthermore, Observation 4.6 implies that $h(x)=4$ for each $x\in U$. Then, $G$ is DP $h$-colorable by Theorem 1.5, a contradiction.
Next, suppose that $\rho_{G^{\prime},h}(U)\leq-2$.
Then $G[U\cup\{v\}]$ is constructed by deleting an edge $uw$ from $G^{\prime}[U]$, adding a vertex $v$ with $h(v)=2$, and adding two edges $vu$ and $vw$; hence,
$\rho_{G,h}(U\cup\{v\})\leq-2+(2\lambda-1)-2(2\lambda)+(2\lambda+1)=-2,$
contradicting Observation 4.6.
Therefore, $G^{\prime\prime}$ is a copy of $K_{k}$, and for some $\varepsilon\in\{0,1\}$, h(x)=k1h(x)=k-1 for $k-1-\varepsilon$ vertices $x\in U$, and h(y)=k2h(y)=k-2 for $\varepsilon$ vertices $y\in U$. Then, writing $U^{\prime}=U\cup\{v\}$,
$\rho_{G,h}(U^{\prime})=k+(2\lambda-1)-2(2\lambda)+2\lambda-\varepsilon(\lambda+2)=k-1-\varepsilon(\lambda+2).$
For $k\geq 5$,
$2\lambda-3=2\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil-3\geq\frac{2k^{2}-14}{2k-7}-3=k-1+\frac{3k}{2k-7}>k-1,$
so by Lemma 4.5, $j:=|E_{G}(U^{\prime},\overline{U^{\prime}})|\leq 1$ and if $\varepsilon=1$, then $j=0$.
If $j=0$, then $U^{\prime}=V(G)$, every vertex of $G$ is low, and $G$ is not a GDP-tree. Therefore, $G$ is DP $h$-colorable, a contradiction. If $j=1$, then $\varepsilon=0$. Write $x$ for the unique vertex of $U^{\prime}$ with a neighbor outside of $U^{\prime}$. As $d(v)=2$, $x\neq v$. If $x\not\in\{u,w\}$, then each vertex $z\in U^{\prime}\setminus\{x\}$ is low, and $G^{\prime}[U^{\prime}\setminus\{x\}]$ is $2$-connected and not a GDP-tree. Hence, $G^{\prime}[U^{\prime}\setminus\{x\}]$ is DP $(h-d_{G}+d_{G[U^{\prime}\setminus\{x\}]})$-colorable, contradicting Observation 4.2.
Otherwise, $x\in\{u,w\}$. We claim that $G[U^{\prime}]$ is DP $(h-d_{G}+d_{G[U^{\prime}]})$-colorable. For each $z\in U^{\prime}$, we write $h^{\prime}(z)=h(z)-d_{G}(z)+d_{G[U^{\prime}]}(z)$, and observe that $h^{\prime}(x)=k-2\geq 3$, and $h^{\prime}(z)=h(z)$ for each $z\in U^{\prime}\setminus\{x\}$. Consider an $h^{\prime}$-cover $(H^{\prime},L^{\prime})$ of $G[U^{\prime}]$. We assign a color $c\in L^{\prime}(x)$ to $x$ with no neighbor in $L(v)$; this is possible, as $|L(v)|=h^{\prime}(v)=2<|L^{\prime}(x)|$. Afterward, each vertex $z\in U^{\prime}\setminus\{x\}$ has at least $d_{G[U^{\prime}\setminus\{x\}]}(z)$ colors in $L^{\prime}(z)$ not adjacent to $c$. Furthermore, $v$ has two colors in $L(v)$ not adjacent to $c$, while $v$ has only one neighbor in $U\setminus\{x\}$. We give a linear ordering to the vertices of $U^{\prime}\setminus\{x\}$ that ends with $v$, and observe that for each $z\in U^{\prime}\setminus\{x\}$, the number of neighbors of $z$ appearing before $z$ in the ordering is less than $|L^{\prime}(z)\setminus\{c\}|$. Therefore, by greedily coloring the vertices of $U^{\prime}\setminus\{x\}$ according to our linear ordering, we complete an $(H^{\prime},L^{\prime})$-coloring of $G[U^{\prime}]$, contradicting Observation 4.2. Hence we can finish an $(H^{\prime},L^{\prime})$-coloring of $G[U^{\prime}]$ using degeneracy, contradicting Observation 4.2. ∎
Corollary 4.12.
Every vertex $v$ in $G$ satisfies $h(v)\geq 3$.
Proof.
This follows from Observation 4.1 and Lemmas 4.9 and 4.11. ∎
Lemma 4.13.
Let $k\geq 5$. If $G$ is $2$-connected and h(v)=k1h(v)=k-1 for each $v\in V(G)$, then $G$ is not (k1)(k-1)-regular.
Proof.
Suppose $G$ is (k1)(k-1)-regular. As $G$ is $h$-minimal, Theorem 1.5 implies that $G$ is either ${\frac{k-1}{t-1}}K_{t}$ for some $2\leq t\leq k-2$ (satisfying (t1)|(k1)(t-1)|(k-1)) or ${(k-1)/2}C_{n}$ for some $n\geq 3$ (provided $k$ is odd). If $G$ is ${\frac{k-1}{t-1}}K_{t}$, then
$\rho_{h}(G)=t((k-1)\lambda+1)-\binom{t}{2}\left(\frac{k-1}{t-1}\right)(2\lambda+1)+\binom{t}{2}=\binom{t+1}{2}-\frac{1}{2}t(k-1)=\frac{1}{2}t(t-k+2).$
As a function of $t$, this expression is a convex quadratic function and is maximized at $t=2$. When $t=2$ and $k\geq 6$, we have $\rho_{h}(G)=2-k+2\leq-2$. Similarly, if $k\geq 6$ and $G={(k-1)/2}C_{n}$ (so $k$ is odd), then
$\rho_{h}(G)$ $=$ $n((k-1)\lambda+1)-n\left(\frac{k-1}{2}\right)(2\lambda+1)+n=n\left(\frac{5}{2}-\frac{1}{2}k\right)\leq-3$
for $k\geq 7$ and $n\geq 3$.
In both cases, $\rho_{h}(G)\leq-2$, so $G$ is not a counterexample to Theorem 2.1, a contradiction.
So suppose $k=5$. If $G=2C_{n}$, then Part 2 of Theorem 2.1 holds. If $G={\frac{5-1}{t-1}}K_{t}$ for some $2\leq t\leq 5-2$, then $t\in\{2,3\}$. If $t=2$, then $G$ is isomorphic to $4K_{2}$. If $t=3$, then $G$ is isomorphic to $2C_{3}$. In all cases, Part 2 of Theorem 2.1 holds. ∎
4.2. A special set $S_{0}^{*}$ of vertices in $G$
An edge-block in a multigraph $G$ is an inclusion maximal connected subgraph $G^{\prime}$ of $G$ such that either $|V(G^{\prime})|=2$ or $G^{\prime}$ has no cut edges. In particular, every cut edge forms an edge-block, and each connected multigraph decomposes into edge-blocks.
Define a special subset $S^{*}_{0}\subseteq V(G)$ as follows. If $G$ has no cut edges, then $S^{*}_{0}=V(G)$. Otherwise, we fix a smallest pendent edge-block $B^{*}$ distinct from $K_{2}$ and let $S^{*}_{0}=V(B^{*})$. By Lemma 4.9 and Observation 4.3, a cut edge cannot be a pendent edge-block, so $S^{*}_{0}$ is well-defined.
If $B^{*}\subsetneq G$, then since $B^{*}$ is pendent, there are $x^{*}_{0}\in S^{*}_{0}$ and $y^{*}_{0}\in V(G)-S^{*}_{0}$ such that $x^{*}_{0}y^{*}_{0}$ is the unique edge connecting $B^{*}$ with the rest of $G$. Fix these $B^{*}$, $S^{*}_{0}$, $x^{*}_{0},y^{*}_{0}$. By definition, $B^{*}$ is $2$-edge-connected.
We have the following lemma.
Lemma 4.14.
Suppose that $S^{*}_{0}\subsetneq V(G)$. For each $S\subseteq V(G)$ satisfying $\emptyset\subsetneq S\cap S_{0}^{*}\subsetneq S_{0}^{*}$, $\rho_{G,h}(S)\geq 2\lambda-3$.
Proof.
If $G[S]$ is not connected, then each vertex set $S^{\prime}$ inducing a component of $G[S]$ is a proper subset of $V(G)$, and hence $\rho_{G,h}(S^{\prime})\geq\lambda-1$ by Lemma 4.5. As $G[S]$ has at least two components, it follows that $\rho_{G,h}(S)\geq 2\lambda-2$. Thus, we assume that $G[S]$ is connected.
Recall that $x_{0}^{*}y_{0}^{*}$ is the cut edge joining $S_{0}^{*}$ and $\overline{S_{0}^{*}}$. First, suppose that $S\subsetneq S_{0}^{*}$. If $\rho_{G,h}(S)\leq 2\lambda-4$, then by Lemma 4.5, $|E_{G}(S,\overline{S})|\leq 1$. Then, $S$ and $\overline{S}\cap S_{0}^{*}$ are joined by at most one edge, contradicting the $2$-edge-connectedness of $B^{*}$.
Next, suppose that $S$ is not contained in $S_{0}^{*}$, so that $x_{0}^{*},y_{0}^{*}\in S$. We write $G^{\prime}=G-S_{0}^{*}$ and observe that $\rho_{G,h}(V(G^{\prime}))\leq\lambda$ by Lemma 4.8. We make several claims.
First, we claim that $\rho_{G,h}(S\cap S_{0}^{*})+\rho_{G,h}(V(G^{\prime}))\geq 4\lambda-3$. Indeed, as $S\cap S_{0}^{*}$ is joined to $S_{0}^{*}-S$ by at least two edges, Lemma 4.5 implies that
$2\lambda-3\leq\rho_{G,h}(S\cup(V(G^{\prime}))\leq-2\lambda+\rho_{G,h}(S\cap S_{0}^{*})+\rho_{G,h}(V(G^{\prime})).$
Rearranging, $\rho_{G,h}(S\cap S_{0}^{*})+\rho_{G,h}(V(G^{\prime}))\geq 4\lambda-3$.
Next, we claim that $\rho_{G,h}(S\cap V(G^{\prime}))\geq\rho_{G,h}(V(G^{\prime}))$. Indeed, if $V(G^{\prime})\subseteq S$, then the statement is clearly true. If $S\cap V(G^{\prime})\subsetneq V(G^{\prime})$, then as $G^{\prime}$ is connected, $E_{G}(S\cap V(G^{\prime}),\overline{S\cap V(G^{\prime})})$ contains $x_{0}^{*}y_{0}^{*}$ as well as at least one additional edge of $G^{\prime}$. Therefore, by Lemma 4.5, $\rho_{G,h}(S\cap V(G^{\prime}))\geq 2\lambda-3>\rho_{G,h}(V(G^{\prime}))$.
Now, putting these claims together,
$\rho_{G,h}(S)=-2\lambda+\rho_{G,h}(S\cap S_{0}^{*})+\rho_{G,h}(S\cap V(G^{\prime}))\geq-2\lambda+\rho_{G,h}(S\cap S_{0}^{*})+\rho_{G,h}(V(G^{\prime}))\geq 2\lambda-3.$
4.3. Subgraphs induced by low vertices in $G$
Let $\Lambda$ be the subgraph of $G$ induced by low vertices in $G$ and let ${\Lambda}_{0}$ be the subgraph of $\Lambda$ induced by the vertices of $\Lambda$ in $S^{*}_{0}$. By Lemma 4.4, each block in $\Lambda$ is a multiple of either a complete graph or a cycle. In particular, each such block is regular. Given a (k2)(k-2)-regular block $B\subseteq\Lambda$ along with distinct vertices $u,u^{\prime}\in N(B)\setminus V(B)$, we write $F(B,u,u^{\prime})$ for the multigraph obtained from GV(B)G-V(B) by adding an edge $uu^{\prime}$.
Lemma 4.15.
If $B$ is a (k2)(k-2)-regular block of $\Lambda$, then the multigraph $F(B,u,u^{\prime})$ is not DP $h$-colorable.
Proof.
As $G$ is $h$-minimal, GV(B)G-V(B) has an $(H,L)$-coloring $\varphi$ that does not extend to all of $G$. For each $v\in V(B)$, we write $L_{\varphi}(v)$ for the subset of $L(v)$ consisting of colors with no neighbor in $\varphi$. We write $H^{\prime}=H[L_{\varphi}(V(B))]$. As $|L_{\varphi}(v)|\geq k-2$ for each $v\in V(B)$ and $B$ has no $(H^{\prime},L_{\varphi})$-coloring, Lemma 3.1 implies that $(H^{\prime},L_{\varphi})$ is one of the blowup covers described in Lemma 3.1. In particular, $H^{\prime}$ is (k2)(k-2)-regular. Thus, $H_{\varphi}:=H[L(V(B))]-V(H^{\prime})$ is isomorphic to a subgraph of $\widetilde{B}$.
Now, we consider two cases. First, suppose that every component of $H[L(V(B))]$ is isomorphic to $H_{\varphi}$. Then, as at least one vertex $v\in V(B)$ satisfies h(v)=k1h(v)=k-1 (namely the neighbor of $u$ in $B$), it follows that $H[L(V(B))]$ is isomorphic to k1k-1 disjoint copies of $H_{\varphi}$, each of which contains exactly one vertex from $L(v)$ for each $v\in V(B)$. Now, we construct an $h$-cover $(H^{\prime\prime},L)$ of $F(B,u,u^{\prime})$ as follows. We begin with $(H,L)$, and we remove $L(V(B))$ from $H$. Then, we add a perfect matching between $L(u)$ and $L(u^{\prime})$ so that $c\in L(u)$ and $c^{\prime}\in L(u^{\prime})$ are adjacent if and only if $c$ and $c^{\prime}$ are adjacent to a common component of $H[L(V(B))]$. Then, if GV(B)G-V(B) has an $(H^{\prime\prime},L)$-coloring $\varphi^{\prime}$, the colors $\varphi^{\prime}(u)$ and $\varphi^{\prime}(u^{\prime})$ are adjacent to distinct components of $H[L(V(B))]$. Hence, there is no $H_{\varphi}$-component in $H[L(V(B))]$ all of whose vertices are adjacent to $\varphi^{\prime}$; therefore, by Lemma 3.1, $\varphi^{\prime}$ can be extended to all of $G$. This contradicts the initial assumption that $G$ is $h$-minimal, completing the first case.
By Lemma 3.1, the only other case to consider is that $H[L(V(B))]$ has a single component isomorphic to $H_{\varphi}$. In this case, $H^{\prime}$ is a $q$-blowup of either a clique or a cycle, where $q\geq 2$ is the quotient of k2k-2 and the regularity of $\widetilde{B}$. In particular, $\widetilde{B}$ has maximum degree less than k2k-2. Then, we construct an $h$-cover $(H^{\prime\prime},L)$ of $F(B,u,u^{\prime})$ as follows. We begin with $(H,L)$, and we remove $L(V(B))$ from $H$. Then, we add a single edge to $H^{\prime\prime}$ joining $\varphi(u)$ and $\varphi(u^{\prime})$.
Now, suppose that $F(B,u,u^{\prime})$ has an $(H^{\prime\prime},L)$-coloring $\varphi^{\prime}$. By construction, either $\varphi^{\prime}(u)\neq\varphi(u)$ or $\varphi^{\prime}(u^{\prime})\neq\varphi(u^{\prime})$. Therefore, some vertex of $H_{\varphi}$ has no neighbor in $\varphi^{\prime}$. For each $v\in V(B)$, we let $L^{\prime\prime}(v)$ be the set of colors $c\in L(v)$ for which $c$ has no neighbor in $\varphi^{\prime}$. Then, $H[L^{\prime\prime}(V(B))]$ is not (k2)(k-2)-regular, so by Lemma 3.1, the coloring $\varphi^{\prime}$ extends to an $(H,L)$-coloring of $G$. This contradicts our initial assumption that $G$ is $h$-minimal.
Therefore, $F(B,u,u^{\prime})$ is not DP $h$-colorable. ∎
By Lemma 4.15, for each (k2)(k-2)-regular subgraph $B\subseteq\Lambda$, and any two distinct vertices $u,u^{\prime}\in N(B)\setminus V(B)$, there is an $h$-minimal multigraph $G(B,u,u^{\prime})$ contained in $F(B,u,u^{\prime})$. As $G$ is $h$-minimal, $G(B,u,u^{\prime})$ is not a subgraph of $G$; therefore, $G(B,u,u^{\prime})$ contains $u$ and $u^{\prime}$.
Lemma 4.16.
If $S^{*}_{0}\neq V(G)$, then the vertex $x^{*}_{0}$ is not in a (k2){(k-2)}-regular block of $\Lambda$.
Proof.
Suppose $x^{*}_{0}$ is in a (k2){(k-2)}-regular block $B$ of $\Lambda$. Since $B$ is $2$-connected, $y^{*}_{0}\notin V(B)$. If there exists a vertex $y_{1}^{*}\in N(V(B))\setminus(V(B)\cup\{y^{*}_{0}\})$, let $G^{\prime}=G(B,y^{*}_{0},y_{1}^{*})$. Recall that $y_{0}^{*},y_{1}^{*}\in V(G^{\prime})$. Since $y^{*}_{0}y_{1}^{*}$ is a cut edge in $G^{\prime}$, it is not part of an exceptional graph. Therefore, $\rho_{G^{\prime},h}(V(G^{\prime}))\leq-2$, and hence $\rho_{G,h}(V(G^{\prime}))\leq 2\lambda-2$. However, the multigraph $G[V(G^{\prime})]$ is disconnected. As $G[S^{*}_{0}]$ is an edge-block of $G$, $|E_{G}(V(G^{\prime})\cap S^{*}_{0},S^{*}_{0}\setminus V(G^{\prime}))|\geq 2$; therefore, by Lemma 4.5, $\rho_{G,h}(V(G^{\prime})\cap S^{*}_{0})\geq 2\lambda-3$. Hence, $\rho_{G,h}(V(G^{\prime})\setminus S^{*}_{0})\leq(2\lambda-2)-(2\lambda-3)<\lambda-1$, contradicting Observation 4.7.
Suppose now that $V(B)$ has no neighbor in $V(G)\setminus(V(B)\cup\{y^{*}_{0}\})$. Then every vertex $u\in V(B)-x_{0}^{*}$ satisfies h(u)=k2h(u)=k-2. Let $b=|V(B)|$, so
$\rho_{G,h}(V(B))=-m(B)+(b-1)((k-2)\lambda-1)+((k-1)\lambda+1)-2\lambda b\left(\frac{k-2}{2}\right)=-m(B)+\lambda+2-b.$
If $b\geq 4,$ then $\rho_{G,h}(V(B))\leq\lambda-2$, contradicting Observation 4.7. Furthermore, if $2\leq b\leq 3,$ then $m(B)\geq 2$, so again $\rho_{G,h}(V(B))\leq\lambda-2$, contradicting Observation 4.7. ∎
Lemma 4.17.
If $k\geq 5$, then ${\Lambda}_{0}$ does not contain $K_{k-1}$ formed by (k1)(k-1)-vertices.
Proof.
Suppose there is a set $C=\{v_{1},v_{2},\ldots,v_{k-1}\}\subseteq S^{*}_{0}$ consisting of low (k1)(k-1)-vertices such that $B:=G[C]=K_{k-1}$. For $i\in[1,k-1]$, let $v^{\prime}_{i}$ be the unique neighbor of $v_{i}$ outside of $C$. Let $C^{\prime}=\{v_{1}^{\prime},v_{2}^{\prime},\ldots,v_{k-1}^{\prime}\}$, so that $v_{i}v_{i}^{\prime}\in E(G)$ for each $i\in[1,k-1]$. Since $G$ has no $K_{k}$, $|C^{\prime}|\geq 2$.
For distinct $v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}$, write $G_{s,t}=G(B,v_{s}^{\prime},v^{\prime}_{t})$. We will simultaneously consider two cases: If $\rho_{h}(G_{s,t})\leq-2$ for some distinct $v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}$, or if $G_{s,t}$ is exceptional for every pair of distinct $v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}$. Let $\varrho(C)=\min\{\rho_{h}(G_{s,t}):v^{\prime}_{s},v^{\prime}_{t}\in C^{\prime}\mbox{ and }v^{\prime}_{s}\neq v^{\prime}_{t}\}$.
For $j\in[1,k-1]$, define the function
[…]
For $j\in[1,k-1]$, let $\pi(j)$ be the minimum of $\rho_{G,h}(P)$ over $P\subseteq V(G)\setminus C$ with $|N(P)\cap C|\geq j$. Let $P_{j}\subseteq V(G)\setminus C$ be such that $|N(P_{j})\cap C|\geq j$ and $\rho_{G,h}(P_{j})=\pi(j)$.
We will prove that $\pi(k-1)\leq g(k-1)$. For this, let $j$ be the largest value for which $\rho_{G,h}(P_{j})\leq g(j)$. If j=k1j=k-1, then we are done, so assume j<k1j<k-1. Then $j\geq 2$ since
[…]
Assume there are distinct $i,i^{\prime}\in[1,k-1]$ such that $v^{\prime}_{i}\in P_{j}$ and $v^{\prime}_{i^{\prime}}\not\in P_{j}$. Since $V(G_{i,i^{\prime}})\cup P_{j}\supseteq P_{j}\cup\{v^{\prime}_{i^{\prime}}\}$, $V(G_{i,i^{\prime}})\cup P_{j}$ has at least $j+1$ neighbors in $C$. Thus, Lemma 2.2 implies that
(16) $\pi(j+1)\leq\rho_{G,h}(P_{j}\cup V(G_{i,i^{\prime}}))\leq g(j)+\rho_{G,h}(V(G_{i,i^{\prime}}))-\rho_{G,h}(P_{j}\cap V(G_{i,i^{\prime}})).$
In Case 2, by Lemma 2.4, $\pi(j+1)\leq g(j)-(k-3)(\lambda-1)+2<g(j+1)$, contradicting the maximality of $j$. In Case 1, $\rho_{h}(G_{i,i^{\prime}})\leq-2$, so as both $P_{j}$ and $V(G_{i,i^{\prime}})$ contain $v^{\prime}_{i}\in S_{0}^{*}$, by Lemma 4.14, $\rho_{G,h}(P_{j}\cap V(G_{i,i^{\prime}}))\geq 2\lambda-3$. Thus, $\pi(j+1)\leq g(j)+(-2+2\lambda+1)-(2\lambda-3)=g(j)+2=g(j+1)$, again contradicting the maximality of $j$. Therefore, $C^{\prime}\subseteq J_{k-1}$, so $\pi(k-1)\leq g(k-1)$.
Since $\rho_{G,h}(C)=(k-1)\left((k-1)\lambda+1\right)-2\lambda\binom{k-1}{2}=(\lambda+1)(k-1)$, we get
$\rho_{G,h}(C\cup P_{C^{\prime}}){\leq}\rho_{G,h}(C)+g(k-1)-2\lambda(k-1){=}(\lambda+1)(k-1)+g(k-1)-2\lambda(k-1){=}(1-\lambda)(k-1)+g(k-1)$
[…]
This contradicts Observation 4.6 in all cases except when $k=5$ and $\varrho(C)=5$. Since $1<\lambda-1$ when $k=5$, Observation 4.7 implies $C\cup P_{C^{\prime}}=V(G)$. In particular, $G[P_{C^{\prime}}]$ is $K_{5}^{-}$ and $G[C^{\prime}]$ is $K_{4}^{-}$ so $G_{s,t}$ is exceptional only for one pair $s,t$. Yet, this is impossible since we are in Case 2. ∎
Lemma 4.18.
${\Lambda}_{0}$ does not contain a (k2){(k-2)}-regular block.
Proof.
Suppose that ${\Lambda}_{0}$ contains $C=\{v_{1},v_{2},\ldots,v_{b}\}$ such that $B:=G[C]$ is a (k2){(k-2)}-regular block. Let $M$ be the set of (k2)(k-2)-vertices in $C$. For all $v_{i}\in C-M$, let $v^{\prime}_{i}$ be the unique neighbor of $v_{i}$ outside of $C$. Since $G$ is $h$-minimal, GCG-C has an $(H,L)$-coloring $\varphi$. We write $H_{\varphi}$ for the set of colors in $H$ above $B$ that are adjacent to $\varphi$. For each $v_{i}\in C$, we write $L_{\varphi}(v_{i})=L(v_{i})\setminus V(H_{\varphi})$. Since $|L_{\varphi}(v)|=d_{B}(v)$ for each $v\in V(B)$ and $B$ is not $(H[L_{\varphi}(C)],L_{\varphi})$-colorable, Theorem 1.5 implies that $B$ is a multiple of either a cycle of a complete graph.
If $B$ is isomorphic to $K_{k-1}$, then by Lemma 3.1, $H[L_{\varphi}(C)]$ consists of k2k-2 copies of the graph $K_{k-1}$. In particular, $H[L_{\varphi}(C)]$ is (k2)(k-2)-regular. By Lemma 4.17, $M$ is nonempty, so $H_{\varphi}$ is a complete graph on at most k2k-2 vertices; in particular, $H_{\varphi}$ is not (k2)(k-2)-regular.
If $B$ is not isomorphic to $K_{k-1}$, then by Lemma 3.1, every component of $H[L_{\varphi}(C)]$ is a (k2)(k-2)-regular $q$-blowup of a cycle or a complete graph for some $q$. More precisely, writing $t$ for the regularity of $\widetilde{B}$, $q=\frac{k-2}{t}$. Therefore, $H_{\varphi}$ is a subgraph of $\widetilde{B}$, and hence each vertex of $H_{\varphi}$ has degree at most t<k2t<k-2. In particular, $H_{\varphi}$ is not (k2)(k-2)-regular.
As $H_{\varphi}$ is a component of $H[L_{\varphi}(C)]$, which is not (k2)(k-2)-regular,
(17) $\forall v_{i}\in C-M$, the color $\varphi(v_{i}^{\prime})\in L(v^{\prime}_{i})$ is adjacent to the unique color $\alpha_{v_{i}}\in L(v_{i})\cap H_{\varphi}$.
If $M=C$, then
$\rho_{h}(M)=-m(B)+b((k-2)\lambda-1)-2\lambda\frac{b(k-2)}{2}=-m(B)+b\big((k-2)\lambda-1-(k-2)\lambda\big)=-m(B)-b\leq-4,$
contradicting Observation 4.6. So $M\neq C$.
Fix a vertex $v_{i}\in C-M$, let $G_{i}=G-C$, and let $L_{i}$ differ from $L|_{V(G)-C}$ only in that $\varphi(v_{i}^{\prime})$ is deleted from $L(v^{\prime}_{i})$. Correspondingly, $h_{i}$ differs from $h$ only in that $h_{i}(v^{\prime}_{i})=h(v^{\prime}_{i})-1$. By (17), $G_{i}$ is not $(H[L_{i}(V(G_{i}))],L_{i})$-colorable. So, there is an $h_{i}$-minimal subgraph $G^{\prime}_{i}$ of $G_{i}$, and it contains $v^{\prime}_{i}$. Since $h_{i}(v^{\prime}_{i})\leq k-2$, $G^{\prime}_{i}\neq K_{k-1}$, and $\rho_{G_{i},h_{i}}(V(G^{\prime}_{i}))\leq-2$. Then $\rho_{G,h}(V(G^{\prime}_{i}))\leq-2+(\lambda+2)=\lambda.$ So, by Lemma 4.5, there is a unique edge joining $V(G^{\prime}_{i})$ with $V(G)\setminus V(G^{\prime}_{i})$. Since $C\cap V(G^{\prime}_{i})=\emptyset$, this edge is $v_{i}v^{\prime}_{i}$. Hence each edge connecting $C$ with V(G)CV(G)-C is a cut edge. As $B^{*}$ is $2$-edge-connected, it follows that $x_{0}^{*}\in C$, contradicting Lemma 4.16. ∎
5. Discharging
We will use the notions of $\sigma_{h}(F)$, $\widetilde{d}(v)$, and $m(F)$ from Lemma 3.2. By Corollary 4.12, each vertex $v\in V(G)$ satisfies $h(v)\geq 3$.
We show that $\rho_{h}(G)\leq-2$, proving that $G$ in fact is not a counterexample to Theorem 2.1. We use the following discharging procedure. We write $\alpha=\frac{k-2}{2k-7}$.
  1. (1)
    For each $v\in V(G)$, we give initial charge $\rho_{h}(v)$ to $v$. For each pair $uv$, where $u,v\in V(G)$ are joined by $s\geq 1$ edges, we give initial charge $s(2\lambda+1)+1$ to the pair $uv$. For each pair $uv$ of non-adjacent $u,v\in V(G)$, the initial charge of $uv$ is 0.
  2. (2)
    For each pair $uv$ of adjacent vertices in $G$, if $s\geq 1$ edges connect $u$ with $v$, the pair $uv$ receives charge $(s(2\lambda+1)-1)/2$ from each of $u$ and $v$.
  3. (3)
    Each non-low vertex $u\in S^{*}_{0}$ takes charge $\alpha$ along each edge $e$ that joins $u$ to a low vertex $v\in S^{*}_{0}$.
Now, we consider the final charges of the vertices in $G$. For each $v\in V(G)$, we write $ch^{*}(v)$ for the final charge of $v$. We observe that the total charge in $G$ is $\rho_{h}(G)$. Additionally, the final charge of each vertex pair is 0, so $\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)$. Finally, if $S^{*}_{0}\neq V(G)$, then the total charge in $G[\overline{S^{*}_{0}}]$ is at most $\rho_{G,h}(\overline{S^{*}_{0}})-\lambda\leq 0$ by Lemma 4.8. Therefore,
(18) $\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)\leq\sum_{v\in S^{*}_{0}}ch^{*}(v).$
For each $v\in S^{*}_{0}$, if $d(v)\geq 1+h(v)$, then we consider several cases.
  1. (N1)
    If $d(v)\leq k-1$, then $h(v)\leq k-2$, so
    $ch^{*}(v)\leq(h(v)\lambda-1)-d(v)\lambda+\alpha d(v)=\lambda(h(v)-d(v))-1+\alpha d(v)\leq-\lambda-1+\alpha d(v)<-1.$
  2. (N2)
    If $d(v)=k$, then
    $ch^{*}(v)\leq((k-1)\lambda+1)-k\lambda+\alpha k=-\lambda+1+\alpha k\leq 0.$
  3. (N3)
    If $d(v)\geq k+1$, then
    $ch^{*}(v)$ $\leq$ $((k-1)\lambda+1)-d(v)\lambda+\alpha d(v)$
    $\leq$ $((k-1)\lambda+1)-(k+1)\lambda+\alpha(k+1)=(-\lambda+1+\alpha k)-\lambda+\alpha<-1.$
For each $v\in S^{*}_{0}$, if $d(v)=h(v)=j$ for some fixed $j\in[3,\ldots,k-1]$, then we consider two cases:
  1. (L1)
    If $j\leq k-2$, then
    $ch^{*}(v)\leq(j\lambda-1)-j\lambda-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\Lambda}_{0}}(v))=-1-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\Lambda}_{0}}(v)).$
  2. (L2)
    If j=k1j=k-1, then
    $ch^{*}(v)$ $\leq$ $(k-1)\lambda+1-(k-1)\lambda-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha(d_{B^{*}}(v)-d_{{\Lambda}_{0}}(v))$
    $=$ $1-\frac{d(v)-\widetilde{d}(v)}{2}-\alpha d_{B^{*}}(v)+\alpha d_{{\Lambda}_{0}}(v).$
We claim that ${\Lambda}_{0}$ is nonempty. Indeed, suppose that ${\Lambda}_{0}$ is empty. As $d(v)\geq 1+h(v)$ for each $v\in S_{0}^{*}$ we have the following inequality instead of (N1)–(N3):
$ch^{*}(v)\leq(h(v)\lambda+1)-d(v)\lambda=(h(v)-d(v))\lambda+1\leq-\lambda+1<-2.$
Then, by (18), $\rho_{h}(G)<-2|S^{*}_{0}|\leq-2$, so $G$ is not a counterexample. Thus, we assume that ${\Lambda}_{0}\neq\emptyset$, or equivalently, that $S_{0}^{*}$ contains at least one low vertex.
Lemma 5.1.
If $B$ is a component of ${\Lambda}_{0}$, then $\sum_{v\in V(B)}ch^{*}(v)<-1$.
Proof.
We note that $d_{B^{*}}(v)=h(v)$ for $v\in V(B)\setminus\{x^{*}_{0}\}$, and $d_{B^{*}}(x^{*}_{0})=h(x^{*}_{0})-1$ whenever $x^{*}_{0}\in V({\Lambda}_{0})$. Hence, by (L1)–(L2),
$\sum_{v\in V(B)}ch^{*}(v)\leq|V_{k-1}(B)|-|V_{k-1}^{-}(B)|-\sum_{v\in V(B)}\left(\frac{d(v)-\widetilde{d}(v)}{2}\right)-\alpha\sum_{v\in V(B)}\left(h(v)-d_{{\Lambda}_{0}}(v)\right)+\alpha$
$|V_{k-1}(B)|-|V_{k-1}^{-}(B)|-m(B)-\alpha\sigma_{h}(B)+\alpha=-\Phi_{k}(B)+\alpha.$
Note that the $\alpha$ term accounts for the possibility that $x^{*}_{0}\in V(B)$, in which case $d_{B^{*}}(x^{*}_{0})=-h(x^{*}_{0})+1$.
By Corollary 4.12, Condition (i) of Lemma 3.2 holds for $T=B$. Condition (ii) holds because the vertices in $B$ are low. By Lemma 4.18, $B$ has no (k2)(k-2)-regular block. If $B$ is (k1)(k-1)-regular, then as $B$ is low and $G$ is connected, $G=B$, which contradicts Lemma 4.13. Therefore, Condition (iii) holds. So, by Lemma 3.2, $\sum_{v\in V(B)}ch^{*}(v)\leq-\Phi_{k}(B)+\alpha<-1$. ∎
Now, by (N1)–(N3), the vertices of $S_{0}^{*}\setminus{\Lambda}_{0}$ have total nonpositive charge. Therefore,
$\rho_{h}(G)=\sum_{v\in V(G)}ch^{*}(v)\leq\sum_{v\in S^{*}_{0}}ch^{*}(v)\leq\sum_{v\in{\Lambda}_{0}}ch^{*}(v).$
As ${\Lambda}_{0}\neq\emptyset$, Lemma 5.1 implies that $\sum_{v\in{\Lambda}_{0}}ch^{*}(v)<-1$. Therefore, $\rho_{h}(G)<-1$, and as $\rho_{h}(G)$ is integral, $\rho_{h}(G)\leq-2$. This completes the proof.
6. Bound for list coloring
We say that $G$ is $(H,L)$-minimal if $G$ has no $(H,L)$-coloring but every proper subgraph $G^{\prime}$ of $G$ has an $(H[L(V(G^{\prime}))],L|_{V(G^{\prime})})$-coloring.
Lemma 6.1.
Let $k\geq 5$, and let $G$ be a multigraph with a function $h:V(G)\to[0,k-1]$. Suppose that $G$ has no exceptional subgraph. If $(H,L)$ is an $h$-cover of $G$ and $G$ is $(H,L)$-minimal, then $\rho_{h}(G)\leq-2$.
Proof.
Suppose for the sake of contradiction that $\rho_{h}(G)\geq-1$. As $G$ is $(H,L)$-minimal, no vertex $v\in V(G)$ satisfies $h(v)=0$. By Theorem 2.1, $G$ is not $h$-minimal. Since $G$ is not DP $h$-colorable, $G$ has a proper subgraph $G_{0}$ that is $h$-minimal. Since $G$ has no exceptional subgraph, Theorem 2.1 implies that $\rho_{G,h}(V(G_{0}))\leq-2$. We let $G_{1}\supseteq G_{0}$ be a maximal subgraph of $G$ for which $\rho_{G,h}(V(G_{1}))\leq-2$, and we observe that $V(G_{1})\subsetneq V(G)$. By the $(H,L)$-minimality of $G$, $G_{1}$ has an $(H,L)$-coloring $\varphi$.
Now, define $G^{\prime}=G-V(G_{1})$. We observe that for each $v\in V(G^{\prime})$, if $|E_{G}(v,V(G_{1}))|=j$, then $h(v)\geq j+1$. Indeed, if $j=0$, then the inequality is clear. If $j\geq 1$ and $h(v)\leq j$, then $\rho_{G,h}(V(G_{1})\cup\{v\})\leq-2+j\lambda+1-2j\lambda<-2$, contradicting the maximality of $G_{1}$.
For each vertex $v\in V(G^{\prime})$, write $h^{\prime}(v)=h(v)-|E_{G}(v,V(G_{1}))|$. By our previous observation, $h^{\prime}(v)\geq 1$ for each $v\in V(G^{\prime})$. Consider a vertex subset $U\subseteq V(G^{\prime})$, and let $j=|E_{G}(U,V(G_{1}))|$. By the maximality of $G_{1}$,
$1\leq\rho_{G,h}(V(G_{1})\cup U)\leq-2-2j\lambda+\rho_{G,h}(U).$
Rearranging, $\rho_{G,h}(U)\geq 1+2j\lambda$, which implies that $\rho_{G,h^{\prime}}(U)\geq 1+j(\lambda-2)$. Thus, every subgraph $G^{\prime\prime}$ of $G^{\prime}$ satisfies $\rho_{h^{\prime}}(G^{\prime\prime})\geq 1$, and hence by Theorem 2.1, $G^{\prime}$ has no $h^{\prime}$-critical subgraph. Therefore, $G^{\prime}$ is DP $h^{\prime}$-colorable, and thus $\varphi$ can be extended to an $(H,L)$-coloring of $G$. This contradicts the original assumption that $G$ is $(H,L)$-minimal, completing the proof. ∎
Now, with Lemma 6.1, we can prove Theorem 1.7.
Theorem 1.7.
If $k\geq 5$, $\lambda=\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil$, and $n\geq k+2$, then $f_{\ell}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}$.
Proof of Theorem 1.7.
Let $G$ be a list $k$-critical multigraph on $n\geq k+2$ vertices and $m$ edges. As $G$ has no $k$-critical proper subgraph, $G$ is $K_{k}$-free. Furthermore, as $G$ is list $k$-critical, $G$ has no parallel edges. Therefore, $G$ has no exceptional subgraph. Let $L^{\prime}:V(G)\to\operatorname{Pow}(\mathbb{N})$ be a list assignment for which $|L^{\prime}(v)|\geq k-1$ for every $v\in V(G)$ and $G$ has no $L^{\prime}$-coloring.
Now, we construct a (k1)(k-1)-cover $(H,L)$ of $G$ for which the $L^{\prime}$-coloring and the $(H,L)$-coloring problems on $G$ are equivalent. As $G$ has no $L^{\prime}$-coloring and is list $k$-critical, it follows that $G$ is $(H,L)$-minimal. Therefore, by Lemma 6.1, $\rho_{h}(G)=(k\lambda+1)n-2\lambda m\leq-2$. Rearranging, $m\geq(k-1+\frac{1}{\lambda})\frac{n}{2}+\frac{1}{\lambda}$. ∎
Concluding remarks. We do not think that our bounds are sharp. On the other hand, it seems that $f_{\mathrm{DP}}(n,k)<f(n,k)$ for many values of $n$ and $k$. An interesting problem is to find the asymptotics of $\frac{f_{\mathrm{DP}}(n,k)}{n}$ for a fixed $k$ and $n\to\infty$.